# Measuring Data Science Automation: A Survey of Evaluation Tools for AI Assistants and Agents

### v1

@@ 3 Evaluating LLM assistants in data science
Considering Data Management activities, Spider (Yu et al., 2018), SParC (Yu et al., 2019b) and CoSQL (Yu et al., 2019a) (all from the same research group) evaluate conversational database querying systems translating natural language into SQL queries (part of Data Architecting but also requiring Data Understanding). These works build on the same 200 databases from 138 domains: Spider consists of 10,181 manually crafted questions and 5,693 unique SQL reference queries and evaluates the generated queries with matching of SQL components or the overall query to the reference one, or with the accuracy of the execution. SParC expands Spider, which contains only single-turn questions, by simulating multi-turn interactions and therefore introducing context dependence: annotators chained Spider tasks together in a conversational flow resulting in 4,298 question sequences with 12,726 questions. Performance is evaluated in terms of exact set match (per turn), and interaction match (full sequence accuracy); however, this does not evaluate the ability of the AI system to interact with a user successfully. This is done in CoSQL, which also includes task where the system must identify ambiguous questions needing clarifications and unanswerable queries (accuracy is evaluated using dialogue act labels). This makes CoSQL unique in addressing intermediate levels of automation for assistants (Sec. 2.2). The clarifications are then included in the context the system uses to determine the correct SQL query, scored using exact match or component match. However, despite being inserted in the context of a conversation, only one system answer at a time is evaluated, therefore still considering the paradigm of substitution (Sec. 2.1). Natural language summaries of the query output produced by the system are also evaluated (with the BLEU score). Overall, CoSQL comprises over 30,000 dialogue turns and 10,000 annotated SQL queries, derived from approximately 3,000 dialogues collected by having users interact with a mock interface controlled by an expert and simulating real-world database query scenarios. Finally, the same authors recently introduced (Lei et al., 2024), Spider 2.0-Lite, consisting of 547 test instructions mapped to 158 real databases hosted on BigQuery, Snowflake and SQLite and solely scored based on execution accuracy, and Spider 2.0-Snow, re-hosting the same 547 questions on Snowflake to spotlight one dialect while keeping identical self-contained evaluation.
@@
Moving to the exploratory aspects of data science, LIDA (Dibia, 2023) introduces a system generating data visualisation and infographics by prompting LLMs in a structured manner to provide a summary of the dataset (Data Understanding), formulate data exploration goals (Data Value Exploration), generate code specifications for the visualisations (Goal Exploration), and generate stylised graphics based on the previous output (Narrative Exploration). This also covers aspects of Data Release as it involves making data accessible through visualisations. The system is accompanied by an evaluation tool, based on 57 datasets sourced from the Vega datasets [^3] https://github.com/vega/vega-datasets repository; two metrics are used: visualisation error rate, computed as the percentage of generated visualisations that result in code compilation errors; and visualisation quality, in which GPT-4 (Achiam et al., 2023) is tasked with assessing the quality of the generated visualisations across 6 dimensions: code accuracy, data transformation, goal compliance, visualisation type, data encoding, and aesthetics.
@@
From this overview, it is evident how nearly all LLM assistant evaluation works focus on code generation, and how there is a concentration on the goal-oriented activities of Data Understanding, Data Preparation and Modelling (Table 2), with lack of evaluation works for exploratory and, even more, data management activities, except for Data Value Understanding, and a few works touching on Narrative Exploration, Goal Exploration, Data Source Exploration, Result Exploration, Data Architecting, and Data Release.
@@ 4 Evaluating LLM agents in data science
Goal-oriented Exploratory Data Management Papers BU DU DP M E Dep GE DSE DVE RE NE PE Acq Sim Arch Rel BLADE (Gu et al., 2024) - ✓✓ ✓✓ ✓✓ - - - - ✓✓ - - - - BiasBenchmark (Li et al., 2025b) - ✓✓ - - - - - - - - - - - - - - CSR-Bench (Xiao et al., 2025) - - ✓✓ - - - - - - - - - - Data-Copilot (Zhang et al., 2024a) - ✓✓ - - - - - - - - - - - DA-Code (Huang et al., 2024b) - ✓✓ ✓✓ ✓✓ - - - - - - - - - - DiscoveryBench (Majumder et al., 2024) - ✓✓ - ✓✓ - - - - DSBench (Jing et al., 2024) - ✓✓ - - - - - - DS-Eval (Zhang et al., 2024b) ✓✓ ✓✓ ✓✓ - ✓✓ - - - - ✓✓ - - - - - IDA-Bench (Li et al., 2025a) - ✓✓ - ✓✓ - - - - - - - - - - InfiAgent-DABench (Hu et al., 2024) - ✓✓ - - - - - - - - - - InsightBench (Sahu et al., 2025) ✓✓ - ✓✓ ✓✓ - - - - ✓✓ MLAgentBench (Huang et al., 2024a) - - ✓✓ - - - - - - - - MLE-Bench (Chan et al., 2024) - ✓✓ - - - - - - - - - MLGym (Nathani et al., 2025) - ✓✓ - - - - - - - - - - RE-Bench (Wijk et al., 2024) - - - ✓✓ - - - - - - - - - - - - ScienceAgentBench (Chen et al., 2024) - ✓✓ ✓✓ ✓✓ - ✓✓ - - ✓✓ - - - - - Spider 2.0 (Lei et al., 2024) - ✓✓ - - - - ✓✓ - - - - - - - ✓✓ - SUPER (Bogin et al., 2024) - - - ✓✓ - - - - - - - - - - Tapilot-Crossing (Li et al., 2024) - ✓✓ ✓✓ - - - - - - - - -
@@ 4.1 Works targeting specific goal-oriented activities
Many works aim to evaluate LLM agents on individual goal-oriented activities of the data science pipeline. Starting from Data Understanding, Li et al. (2025b) introduce BiasBenchmark, which evaluates the ability to detect biases in datasets. To build the benchmark, they select 5 datasets from prior bias mitigation research and 100 demographic-related features or their combinations. They then craft (using an LLM playing the role of a user) possible bias detection queries, including intentionally ambiguous questions. During evaluation, clarification questions posed by the agent are automatically answered by the LLM-based user simulator according to the original task specifications, ensuring consistency and reproducibility. The agent must quantify the bias level according to a 5-level scale, which is then compared to the ground truth obtained by measuring five widely-used bias detection metrics. However, scoring the final bias level allows to agent to determine it in potentially novel ways, therefore going above human substitution (Sec. 2.1). They also evaluate the agent’s intermediate process, by developing an agent-based automated evaluation system that looks at the evaluated agent’s logs and produces performance rating levels for five aspects: user communication, task planning, tool invocation, dynamic plan adjustment and result analysis. Interestingly, the first one (user communication) scores the ability of the agent to ask clarification questions to the (simulated) user who set the task.
@@
Instead, Data-Copilot (Zhang et al., 2024a) introduces an LLM agent for data wrangling that, given a dataset schema, independently explores potential user requests and generates modular code to address them, which is then leveraged in the deployment stage. To benchmark it, the authors release 547 test requests drawn from 173 human seeds plus a larger 3547-request self-exploration pool. The tasks rely on financial data and touch upon Data Value Exploration, Data Understanding and Preparation, and Modelling. Each test case comes with a human-curated answer table and four manual labels for dataset analysis: task difficulty, request rationality, expression ambiguity, answer accuracy. System performance is measured with GPT-4-based Pass@1 scoring against the gold tables (plus an image check) and with the number of tokens used.
@@
Moving towards Modelling, MLE-Bench (Chan et al., 2024) and RE-Bench (Wijk et al., 2024) are both learning engineering benchmarks, but differ in the complexity of the tasks and scenarios: MLE-Bench encompasses 75 tasks sourced from Kaggle [^4] https://www.kaggle.com/, whose deterministic scoring functions are taken from the corresponding Kaggle competitions—as they score the final result, this allows agents to solve the task in ways potentially different from humans, going beyond substitution (Sec. 2.1); however, because these functions vary across tasks, each score is compared against a snapshot of the (human) leaderboard. Instead, RE-Bench includes 7 environments each presenting a unique Machine Learning (ML) task focused on optimising either the loss function or the run-time; the value of these scoring functions is manually inspected for evaluation, and evaluators need to have access to a reference solution. Relatedly, MLAgentBench (Huang et al., 2024a) comprises 13 tasks specified by a goal, occasionally constraints or specific instructions, starter files, and an evaluator; the tasks are collected and adapted from recent Kaggle challenges, CLRS (Veličković et al., 2022), BabyLM [^5] https://babylm.github.io/; the starter files consist of data, description of data and metric, and initial code; each task has its own goal metric to improve on, whose measure is used for automated evaluation. For an overview of MLAgent workflow and evaluation, see Fig.  6. Instead, Nathani et al. (2025) introduce MLGym, an environment to train LLM agents on ML tasks using reinforcement learning. Given a task description, an initial codebase, and actions and observations history, the agent generates an action (shell commands executed by the environment) to accomplish research objectives iteratively; the execution feedback can then be used to refine the agent. MLGym is equipped with a benchmark consisting of 13 tasks spanning data science, game theory, computer vision, NLP, and reinforcement learning, and selected from sources such as Kaggle’s House Price Prediction [^6] https://www.kaggle.com/datasets/zafarali27/house-price-prediction-dataset, 3-SAT (Cook, 1971), CIFAR-10’s image classification [^7] https://www.kaggle.com/code/faressayah/cifar-10-images-classification-using-cnns-88, and more; they require the agent to perform Data Understanding, Modelling and Evaluation. As the various tasks have different performance metric, they score each agent by a quantity that reflects how closely, on average across a range of tolerance levels, it matches the best performer on every individual task. MLGym differs from MLAgentBench for the larger complexity of its tasks. Finally, (Li et al., 2025a) introduce IDA-Bench, which attempts to evaluate LLMs on their ability to perform guided predictive Modelling tasks; the benchmark includes an LLM-simulated user with domain knowledge and subjective insights who interacts with an agent to provide instructions throughout a multi-turn iterative process; the agent is then tested on adapting its goal and following instructions. The tasks are obtained from Kaggle; an LLM distils reference insights in natural language format from an optimal solution. These, together with information such as hyperparameters, serve as a task-specific template for the simulated user, which requests the agent to perform certain steps, without offering all insights up-front, and offers clarifications when asked. Results of the trained model are evaluated against a ground truth using task-specific evaluation functions; and compared with a human baseline, obtained by running the notebook the simulated user has access to. They also determine the ability to interact by considering how the prediction accuracy changes by increasing the number of interactions. Fig. 3 shows an example “trajectory” of the guided data analysis process.
@@ 4.2 Evaluating multiple activities explicitly
Moving to a broader range of activities, DSEval (Zhang et al., 2024b) contains chains of inter-dependent problems (based on data from StackOverflow, Pandas-exercises [^10] https://github.com/guipsamora/pandas_exercises, LeetCode [^11] https://leetcode.com/, and Kaggle) where each highlights a different stage of the data-science lifecycle—Data Understanding and Preparation, Modelling, or interpretation (belonging to Deployment and Narrative Exploration)—while re-using the runtime context left by the previous problems. By doing so, agents must solve the overall task by following the same steps that humans would follow; thus, DSEval only evaluates agents’ substitution ability rather than their potential to transform tasks (Sec. 2.1). For each problem, they employ custom validator modules to check correctness against the solution or run unit tests. Relatedly, Tapilot-Crossing (Li et al., 2024) obtains a set of tasks by simulating (with LLMs) a company setup composed of an administrator and data scientist solving a client’s problem making use of an AI Chatbot Agent; they then manually filter those interactions where the Chatbot Agent produced correct code and obtain 1024 interactions where agents are asked to write code to solve a problem or answer a multiple-choice question. Overall, these tasks cover Data Understanding and Preparation, Data Value Exploration, Modelling, Deployment, Results and Narrative Exploration (by converting plots into answers or summarising findings in prose). They test agents both in a “normal” mode, where all requirements and details are specified by the user, and in an “action” mode, where the agent has to perform an action such as asking for clarification, updating code based on user-reported error, and others. For code updating and normal turns, the agent’s interactivity is disabled, effectively falling back to an assistant setup, while in the other cases the agent may iteratively call a sandboxed Python executor. Further, for the “clarification” tasks, the agent has to pose follow-up questions that are answered by a LLM-simulated user, making this one of a few benchmarks (Li et al., 2025b; a) that evaluate interactivity. An example of the different interaction modes can be found in Fig.  4. Overall, however, scoring is purely based on the outcome (thus not judging the quality of the interaction): the agent’s artefact is compared to a gold reference with task-specific comparators. Finally, ScienceAgentBench (Chen et al., 2024) builds on 102 tasks from scientific peer-reviewed publications, validated by subject experts; each task includes a data-driven discovery goal, information on the data, expert-provided knowledge, and a reference Python program. The questions are challenging, such as “Develop a drug-target interaction model with the DAVIS dataset to repurpose the antiviral drugs for COVID”, or “Analyze Toronto fire stations and their service coverage to identify coverage gaps”. The performance of the LLM agent on each task is scored against 3 metrics: Program Evaluation (itself consisting of: Valid Execution Rate, Success Rate, API Cost and embedding similarity computed by CodeBERT Zhou et al., 2023); Figure Evaluation (using GPT-4o); and Rubric-Based Evaluation based on 5 fundamental steps (Data Loading, Data Processing, Modelling or Visualisation, Output Formatting, and Output Saving). Therefore, this mostly evaluates substitution (Sec. 2.1) both by closely referring to human-provided solutions and by splitting the task in the same sequence of steps humans would follow.

### v2

@@ 3 Evaluating LLM assistants in data science
Considering Data Management activities, Spider (Yu et al., 2018), SParC (Yu et al., 2019b) and CoSQL (Yu et al., 2019a) (all from the same research group) evaluate conversational database querying systems translating natural language into SQL queries (part of Data Architecting but also requiring Data Understanding). These works build on the same 200 databases from 138 domains: Spider consists of 10,181 manually crafted questions and 5,693 unique SQL reference queries and evaluates the generated queries with matching of SQL components or the overall query to the reference one, or with the accuracy of the execution. SParC expands Spider, which contains only single-turn questions, by simulating multi-turn interactions and therefore introducing context dependence: annotators chained Spider tasks together in a conversational flow resulting in 4,298 question sequences with 12,726 questions. Performance is evaluated in terms of exact set match (per turn), and interaction match (full sequence accuracy); however, this does not evaluate the ability of the AI system to interact with a user successfully. This is done in CoSQL, which also includes task where the system must identify ambiguous questions needing clarifications and unanswerable queries (accuracy is evaluated using dialogue act labels). This makes CoSQL unique in addressing intermediate levels of automation for assistants (Sec. 2.2). The clarifications are then included in the context the system uses to determine the correct SQL query, scored using exact match or component match. However, despite being inserted in the context of a conversation, only one system answer at a time is evaluated, therefore still considering the paradigm of substitution (Sec. 2.1). Natural language summaries of the query output produced by the system are also evaluated (with the BLEU score). Overall, CoSQL comprises over 30,000 dialogue turns and 10,000 annotated SQL queries, derived from approximately 3,000 dialogues collected by having users interact with a mock interface controlled by an expert and simulating real-world database query scenarios. Finally, the same authors recently introduced (Lei et al., 2024), Spider 2.0-Lite, consisting of 547 test instructions mapped to 158 real databases hosted on BigQuery, Snowflake and SQLite and solely scored based on execution accuracy, and Spider 2.0-Snow, re‑hosting the same 547 questions on Snowflake to spotlight one dialect while keeping identical self‑contained evaluation.
@@
Moving to the exploratory aspects of data science, LIDA (Dibia, 2023) introduces a system generating data visualisation and infographics by prompting LLMs in a structured manner to provide a summary of the dataset (Data Understanding), formulate data exploration goals (Data Value Exploration), generate code specifications for the visualisations (Goal Exploration), and generate stylised graphics based on the previous output (Narrative Exploration). This also covers aspects of Data Release as it involves making data accessible through visualisations. The system is accompanied by an evaluation tool, based on 57 datasets sourced from the Vega datasets [^3] https://github.com/vega/vega-datasets repository; two metrics are used: visualisation error rate, computed as the percentage of generated visualisations that result in code compilation errors; and visualisation quality, in which GPT-4 (Achiam et al., 2023) is tasked with assessing the quality of the generated visualisations across 6 dimensions: code accuracy, data transformation, goal compliance, visualisation type, data encoding, and aesthetics. Instead, AssistedDS (Luo et al. (2025)) is a benchmark designed to evaluate how well LLMs leverage domain knowledge to improve predictive performance on tabular datasets. It tests the model’s ability to critically assess, filter, and apply both helpful and harmful external information, using synthetic datasets with controlled feature-label relationships and Kaggle datasets augmented with high- and low-rated notebook insights. Evaluation focuses on code quality and predictive accuracy across, and the same task is evaluated by providing different configurations of domain knowledge. By measuring how effectively models identify and apply valuable domain insights, AssistedDS directly targets the core of Data Value Exploration.
@@
From this overview, it is evident how nearly all LLM assistant evaluation works focus on code generation, and how there is a concentration on the goal-oriented activities of Data Understanding, Data Preparation and Modelling (Table 2), with a paucity of evaluation works for exploratory and, even more, data management activities, except for Data Value Understanding, and a few works touching on Narrative Exploration, Goal Exploration, Data Source Exploration, Result Exploration, Data Architecting, and Data Release.
@@ 4 Evaluating LLM agents in data science
Goal-oriented Exploratory Data Management Papers BU DU DP M E Dep GE DSE DVE RE NE PE Acq Sim Arch Rel BLADE (Gu et al., 2024) - ✓✓ ✓✓ ✓✓ - - - - ✓✓ - - - - BiasBenchmark (Li et al., 2025b) - ✓✓ - - - - - - - - - - - - - - CoTa (Li et al., 2025c) - ✓✓ ✓✓ - - - - - - - - - CSR-Bench (Xiao et al., 2025) - - ✓✓ - - - - - - - - - - Data-Copilot (Zhang et al., 2024a) - ✓✓ - - - - - - - - - - - DA-Code (Huang et al., 2024b) - ✓✓ ✓✓ ✓✓ - - - - - - - - - - DiscoveryBench (Majumder et al., 2024) - ✓✓ - ✓✓ - - - - DSBench (Jing et al., 2024) - ✓✓ - - - - - - DS-Eval (Zhang et al., 2024b) ✓✓ ✓✓ ✓✓ - ✓✓ - - - - ✓✓ - - - - - IDA-Bench (Li et al., 2025a) - ✓✓ - ✓✓ - - - - - - - - - - InfiAgent-DABench (Hu et al., 2024) - ✓✓ - - - - - - - - - - InsightBench (Sahu et al., 2025) ✓✓ - ✓✓ ✓✓ - - - - ✓✓ MLAgentBench (Huang et al., 2024a) - - ✓✓ - - - - - - - - MLE-Bench (Chan et al., 2024) - ✓✓ - - - - - - - - - MLGym (Nathani et al., 2025) - ✓✓ - - - - - - - - - - RE-Bench (Wijk et al., 2024) - - - ✓✓ - - - - - - - - - - - - ScienceAgentBench (Chen et al., 2024) - ✓✓ ✓✓ ✓✓ - ✓✓ - - ✓✓ - - - - - Spider 2.0 (Lei et al., 2024) - ✓✓ - - - - ✓✓ - - - - - - - ✓✓ - SUPER (Bogin et al., 2024) - - - ✓✓ - - - - - - - - - - WebDS (Hsu et al., 2025) ✓✓ - ✓✓ - - ✓✓ ✓✓ - - - -
@@ 4.1 Works targeting specific goal-oriented activities
Many works aim to evaluate LLM agents on individual goal-oriented activities of the data science pipeline. Starting from Data Understanding, Li et al. (2025b) introduce BiasBenchmark, which evaluates the ability to detect biases in datasets. To build the benchmark, they select 5 datasets from prior bias mitigation research and 100 demographic-related features or their combinations and craft (using an LLM playing the role of a user) possible bias detection queries, including intentionally ambiguous questions. Then, during evaluation, a bias detection query is passed to the evaluated LLM agent to test whether it is able to effectively detect if the bias exists by analysing the provided data; clarification questions posed by the agent are automatically answered by the LLM-based user simulator according to the original task specifications, ensuring consistency and reproducibility. The agent must quantify the bias level according to a 5-level scale, which is then compared to the ground truth obtained by measuring five widely-used bias detection metrics. However, scoring the final bias level allows to agent to determine it in potentially novel ways, therefore going above human substitution (Sec. 2.1). They also evaluate the agent’s intermediate process, by developing an agent-based automated evaluation system that looks at the evaluated agent’s logs and produces performance rating levels for five aspects: user communication, task planning, tool invocation, dynamic plan adjustment and result analysis. Interestingly, the first one (user communication) scores the ability of the agent to ask clarification questions to the (simulated) user who set the task.
@@
Instead, Data-Copilot (Zhang et al., 2024a) introduces an LLM agent for data wrangling that, given a dataset schema, independently explores potential user requests and generates modular code to address them, which is then leveraged in the deployment stage. To benchmark it, the authors release 547 test requests drawn from 173 human seeds plus a larger 3547-request self-exploration pool. The tasks rely on financial data and touch upon Data Value Exploration, Data Understanding and Preparation, and Modelling. Each test case is accompanied by a human-curated answer table and they are jointly (manually) annotated with four labels for dataset analysis: task difficulty, request rationality, expression ambiguity, answer accuracy. System performance is measured with GPT‑4‑based Pass@1 scoring against the gold tables (plus an image check) and with the number of tokens used.
@@
Moving towards Modelling, MLE-Bench (Chan et al., 2024) and RE-Bench (Wijk et al., 2024) are both learning engineering benchmarks, but differ in the complexity of the tasks and scenarios: MLE-Bench encompasses 75 tasks sourced from Kaggle [^4] https://www.kaggle.com/, whose deterministic scoring functions are taken from the corresponding Kaggle competitions—as they score the final result, this allows agents to solve the task in ways potentially different from humans, going beyond substitution (Sec. 2.1); however, because these functions vary across tasks, each score is compared against a snapshot of the (human) leaderboard. Instead, RE-Bench includes 7 environments each presenting a unique Machine Learning (ML) task focused on optimising either the loss function or the run-time; the value of these scoring functions is manually inspected for evaluation, and evaluators need to have access to a reference solution. Relatedly, MLAgentBench (Huang et al., 2024a) comprises 13 tasks specified by a goal, occasionally constraints or specific instructions, starter files, and an evaluator; the tasks are collected and adapted from recent Kaggle challenges, CLRS (Veličković et al., 2022), BabyLM [^5] https://babylm.github.io/; the starter files consist of data, description of data and metric, and initial code; each task has its own goal metric to improve on, whose measure is used for automated evaluation. For an overview of MLAgent workflow and evaluation, see Fig. 7. Instead, Nathani et al. (2025) introduce MLGym, an environment to train LLM agents on ML tasks using reinforcement learning. Given a task description, an initial codebase, and actions and observations history, the agent generates an action (shell commands executed by the environment) to accomplish research objectives iteratively; the execution feedback can then be used to refine the agent. MLGym is equipped with a benchmark consisting of 13 tasks spanning data science, game theory, computer vision, NLP, and reinforcement learning, and selected from sources such as Kaggle’s House Price Prediction [^6] https://www.kaggle.com/datasets/zafarali27/house-price-prediction-dataset, 3-SAT (Cook, 1971), CIFAR-10’s image classification [^7] https://www.kaggle.com/code/faressayah/cifar-10-images-classification-using-cnns-88, and more; they require the agent to perform Data Understanding, Modelling and Evaluation. As the various tasks have different performance metric, they score each agent by a quantity that reflects how closely, on average across a range of tolerance levels, it matches the best performer on every individual task. MLGym differs from MLAgentBench for the larger complexity of its tasks. Finally, (Li et al., 2025a) introduce IDA-Bench, which attempts to evaluate LLMs on their ability to perform guided predictive Modelling tasks; the benchmark includes an LLM-simulated user with domain knowledge and subjective insights who interacts with an agent to provide instructions throughout a multi-turn iterative process; the agent is then tested on adapting its goal and following instructions. The tasks are obtained from Kaggle; an LLM distils reference insights in natural language format from an optimal solution. These, together with information such as hyperparameters, serve as a task-specific template for the simulated user, which requests the agent to perform certain steps, without offering all insights up-front, and offers clarifications when asked. Results of the trained model are evaluated against a ground truth using task-specific evaluation functions; and compared with a human baseline, obtained by running the notebook the simulated user has access to. They also determine the ability to interact by considering how the prediction accuracy changes by increasing the number of interactions. Fig. 4 shows an example “trajectory” of the guided data analysis process.
@@ 4.2 Evaluating multiple activities explicitly
Moving to a broader range of activities, DSEval (Zhang et al., 2024b) contains chains of inter-dependent problems (based on data from StackOverflow, Pandas-exercises [^10] https://github.com/guipsamora/pandas_exercises, LeetCode [^11] https://leetcode.com/, and Kaggle) where each highlights a different stage of the data-science lifecycle—Data Understanding and Preparation, Modelling, or interpretation (belonging to Deployment and Narrative Exploration)—while re-using the runtime context left by the previous problems. By doing so, agents must solve the overall task by following the same steps that humans would follow; thus, DSEval only evaluates agents’ substitution ability rather than their potential to transform tasks (Sec. 2.1). For each problem, they employ custom validator modules to check correctness against the solution or run unit tests. Relatedly, CoTa (Li et al., 2025c) obtains a set of tasks by simulating (with LLMs) a company setup composed of an administrator and data scientist solving a client’s problem making use of an AI Chatbot Agent; they then manually filter those interactions where the Chatbot Agent produced correct code and obtain 1024 interactions where agents are asked to write code to solve a problem or answer a multiple-choice question. Overall, these tasks cover Data Understanding and Preparation, Data Value Exploration, Modelling, Deployment, Results and Narrative Exploration (by converting plots into answers or summarising findings in prose). They test agents both in a “normal” mode, where all requirements and details are specified by the user, and in an “action” mode, where the agent has to perform an action such as asking for clarification, updating code based on user-reported error, and others. For code updating and normal turns, the agent’s interactivity is disabled, effectively falling back to an assistant setup, while in the other cases the agent may iteratively call a sandboxed Python executor. Further, for the “clarification” tasks, the agent has to pose follow-up questions that are answered by a LLM-simulated user, making this one of a few benchmarks (Li et al., 2025b; a) that evaluate interactivity. An example of the different interaction modes can be found in Fig. 5. Overall, however, scoring is purely based on the outcome (thus not judging the quality of the interaction): the agent’s artefact is compared to a gold reference with task-specific comparators. Finally, ScienceAgentBench (Chen et al., 2024) builds on 102 tasks from scientific peer-reviewed publications, validated by subject experts; each task includes a data-driven discovery goal, information on the data, expert-provided knowledge, and a reference Python program. The questions are challenging, such as “Develop a drug-target interaction model with the DAVIS dataset to repurpose the antiviral drugs for COVID”, or “Analyze Toronto fire stations and their service coverage to identify coverage gaps”. The performance of the LLM agent on each task is scored against 3 metrics: Program Evaluation (itself consisting of: Valid Execution Rate, Success Rate, API Cost and embedding similarity computed by CodeBERT Zhou et al., 2023); Figure Evaluation (using GPT-4o); and Rubric-Based Evaluation based on 5 fundamental steps (Data Loading, Data Processing, Modelling or Visualisation, Output Formatting, and Output Saving). Therefore, this mostly evaluates substitution (Sec. 2.1) both by closely referring to human-provided solutions and by splitting the task in the same sequence of steps humans would follow.